1. Problem 8.01 (Expected Return) eBook A stocks returns have the following distribution: Demand for the Companys Products
Answer #1

Expected Return = (0.1(-0.28) + 0.1(-0.13) + 0.5(0.12) + 0.1(0.29) + 0.2(0.63)

Expected Return = 0.174

Standard Deviation = [(0.1(-0.28 - 0.174)2 + 0.1(-0.13 - 0.174)2 + 0.5(0.12 - 0.174)2 + 0.1(0.29 - 0.174)2 + 0.2(0.63 - 0.174)2)]

Standard Deviation = 0.2725

CV = 0.2725/0.174 = 1.57

Sharpe Ratio = (0.174 - 0.02)/0.2725

Sharpe Ratio = 0.57